# Edinburgh, EH9 3FB, UK https://andrewwango.github.io/perspective-equivariant-imaging

### v1

@@ 4.2 Multispectral satellite image pansharpening
Method PSNR $\uparrow$ ERGAS $\downarrow$ Baseline (Brovey) 19.21 6.65 Baseline (linear) 7.01 24.98 (Oracle supervised) (25.22) (4.24) Self-sup (Wald) 20.97 6.33 Competitor: GDD $<0$ $\sim 10^{2}$ Competitor: SSQ $<0$ 93.65 Competitor: Z-PNN $<0$ $\sim 10^{4}$ Competitor: PanGan 17.06 10.27 MC (SURE) 20.52 6.51 Shift EI 22.58 5.34 Affine EI (ours) 21.95 5.61 Perspective-EI (ours) 21.92 6.3 Pan+tilt EI (ours) 23.17 5.02
@@
We observe better results when using our pan+tilt transform compared to the linear shift transform from [13]. However, by compounding transforms into ours from Tab. 1 (affine, full perspective) we observe even better results as we are increasing the structure in the model.
All of the above are exacerbated in the noisy scenario, where the unsupervised competitors degrade in performance, even in our moderate noise scenario. The classical baseline, Brovey, is not competitive in this scenario. Again, Wald’s protocol contains spectral inconsistencies. Perspective EI remains robust to noise, with pan+tilt EI (which has a perspective effect) approaching the oracle supervised performance, which has access to clean GT.
@@ 6.0.1 Proof of Corollary 1
When two different arbitrary cameras at different locations image a planar scene, their image points are related by a projective transformation, since the intersection of rays with the planar scene are uniquely defined [25, Sec. 13]. Points $\mathbf{\bar{x}}$ in the first camera’s image plane can be related to points $\mathbf{\bar{x}}_{\pi}$ on the scene plane $\pi$ via a homography $\mathbf{H}_{1\pi}$ and similarly for the second camera $\mathbf{H}_{2\pi}$, so the transformation from one image plane to another can be written $\mathbf{\bar{x}}^{\prime}=\mathbf{H}_{2\pi}\mathbf{H}_{1\pi}^{-1}\mathbf{\bar{x}}=\mathbf{T}_{g}\mathbf{\bar{x}}$blacksquare
@@ 6.0.2 Proof of Theorem 1
Recall that, in a pinhole camera, any image point $\mathbf{\bar{x}}=\mathbf{K}\left[\mathbf{R}\mid\mathbf{t}\right]\mathbf{\bar{X}}$ where $\mathbf{\bar{X}}$ is the location of the 3D world point $\mathbf{X}$ in 4-vector homogeneous coordinates and $\mathbf{R},\mathbf{t}$ are the orientation and location of the camera coordinate frame wrt. the world coordinate frame [25, Sec. 6.1]. Suppose another arbitrary camera images the same world point $\mathbf{\bar{x}}^{\prime}=\mathbf{K}^{\prime}\left[\mathbf{R}^{\prime}\mid\mathbf{t}^{\prime}\right]\mathbf{\bar{X}}$. Since we have no awareness of absolute world position, assume w.l.o.g. that $\mathbf{R}=\mathbf{I},\mathbf{t}=\mathbf{0}$ i.e. the first camera’s coordinate frame coincides with the world’s. Furthermore, under Proposition 1, $\mathbf{t}^{\prime}=\mathbf{t}$. Therefore $\mathbf{\bar{x}}^{\prime}=\mathbf{K^{\prime}}\mathbf{R}^{\prime}\mathbf{K}^{-1}\mathbf{\bar{x}}$blacksquare
@@ 6.0.4 Derivation of pan and tilt transformation
A perspective effect is when parallel lines are not preserved after transformation. In other words, perspective transformations $\mathbf{T}$ are those that transform points at infinity (i.e. where parallel lines meet) to finite "vanishing" points. Points at infinity can be written in homogeneous coordinates as $(u,v,0)^{\top}$ [25, Sec. 2.2.2]. The transformation can be written most generally using 8 degrees of freedom (because homography is defined up to a scale constant) as:
@@ 6.0.5 Proof of Theorem 2
From [25, Sec. 2.2.2]: the projective space $\mathbb{P}^{2}$ is the set of all homogeneous 3-vectors. This consists of finite points $(u,v,1)^{\top}\in\mathbb{R}^{2}$ endowed with the set of "points at infinity" $(u,v,0)^{\top}$ which all lie on the line at infinity.

### v2

@@ 4.2 Multispectral satellite image pansharpening
Method PSNR $\uparrow$ ERGAS $\downarrow$ SAM $\downarrow$ Baseline (Brovey) 19.21 6.65 0. Baseline (linear) 7.01 24.98 0.999 (Oracle supervised) (25.22) (4.24) (0.116) Self-sup (Wald) 20.97 6.33 0.156 Competitor: GDD $<0$ $\sim 10^{2}$ 1.461 Competitor: SSQ $<0$ 93.65 1.174 Competitor: Z-PNN $<0$ $\sim 10^{4}$ 1.540 Competitor: PanGan 17.06 10.27 0.206 MC (SURE) 20.52 6.51 0.194 Shift EI 22.58 5.34 0.144 Affine EI (ours) 21.95 5.61 0.151 Perspective-EI (ours) 21.92 6.3 0.155 Pan+tilt EI (ours) 23.17 5.02 0.140
@@
We observe better results when using all our proposed transforms from Tab. 1 compared to the linear shift transform from [13], as we are increasing the structure in the model. Transforms with a perspective or pan+tilt effect perform the best. As perspective-EI’s no-reference quality is the best, we suspect that the reference metrics break down as the quality surpasses that of the reference.
All of the above are exacerbated in the noisy scenario, where the unsupervised competitors degrade in performance, even in our moderate noise scenario. The classical baseline, Brovey, is not competitive in this scenario. Again, Wald’s protocol contains spectral inconsistencies. Perspective-EI remains robust to noise, with pan+tilt EI approaching the oracle supervised performance (idem above for comparison), which has access to clean GT (impossible to obtain in real-world).
@@ 6.0.1 Proof of Corollary 1
When two different arbitrary cameras at different locations image a planar scene, their image points are related by a projective transformation, since the intersection of rays with the planar scene are uniquely defined [26, Sec. 13]. Points $\mathbf{\bar{x}}$ in the first camera’s image plane can be related to points $\mathbf{\bar{x}}_{\pi}$ on the scene plane $\pi$ via a homography $\mathbf{H}_{1\pi}$ and similarly for the second camera $\mathbf{H}_{2\pi}$, so the transformation from one image plane to another can be written $\mathbf{\bar{x}}^{\prime}=\mathbf{H}_{2\pi}\mathbf{H}_{1\pi}^{-1}\mathbf{\bar{x}}=\mathbf{T}_{g}\mathbf{\bar{x}}$blacksquare
@@ 6.0.2 Proof of Theorem 1
Recall that, in a pinhole camera, any image point $\mathbf{\bar{x}}=\mathbf{K}\left[\mathbf{R}\mid\mathbf{t}\right]\mathbf{\bar{X}}$ where $\mathbf{\bar{X}}$ is the location of the 3D world point $\mathbf{X}$ in 4-vector homogeneous coordinates and $\mathbf{R},\mathbf{t}$ are the orientation and location of the camera coordinate frame wrt. the world coordinate frame [26, Sec. 6.1]. Suppose another arbitrary camera images the same world point $\mathbf{\bar{x}}^{\prime}=\mathbf{K}^{\prime}\left[\mathbf{R}^{\prime}\mid\mathbf{t}^{\prime}\right]\mathbf{\bar{X}}$. Since we have no awareness of absolute world position, assume w.l.o.g. that $\mathbf{R}=\mathbf{I},\mathbf{t}=\mathbf{0}$ i.e. the first camera’s coordinate frame coincides with the world’s. Furthermore, under Proposition 1, $\mathbf{t}^{\prime}=\mathbf{t}$. Therefore $\mathbf{\bar{x}}^{\prime}=\mathbf{K^{\prime}}\mathbf{R}^{\prime}\mathbf{K}^{-1}\mathbf{\bar{x}}$blacksquare
@@ 6.0.4 Derivation of pan and tilt transformation
A perspective effect is when parallel lines are not preserved after transformation. In other words, perspective transformations $\mathbf{T}$ are those that transform points at infinity (i.e. where parallel lines meet) to finite "vanishing" points. Points at infinity can be written in homogeneous coordinates as $(u,v,0)^{\top}$ [26, Sec. 2.2.2]. The transformation can be written most generally using 8 degrees of freedom (because homography is defined up to a scale constant) as:
@@ 6.0.5 Proof of Theorem 2
From [26, Sec. 2.2.2]: the projective space $\mathbb{P}^{2}$ is the set of all homogeneous 3-vectors. This consists of finite points $(u,v,1)^{\top}\in\mathbb{R}^{2}$ endowed with the set of "points at infinity" $(u,v,0)^{\top}$ which all lie on the line at infinity.